Benzene-1,2,3-tricarboxylic acid (C6H3 (CO2H) 3) is a three-protic, water-soluble carboxylic acid with the following pKa values: pKa1 = 2.86, pka2 = 4.30 and pka3 = 6.28. What will be the pH of a 0.075 M solution of this treprotic carboxylic acid?
Answer #1


pkai -2• $6 -1°јка, =) а, а löpka, О-оо138 4 = а, - «Нұ(сон, ана (он), coe+ ht I 0.075M Х E c.o1S-X х 14, - [C6H3(CO2H)2(00] x² = 0.00138 %0.075 -0.00138% 2²+0.001384 0.000103550 -0.00138 + (0.00138)2-40)6-0.0001035) 2(1) -0.00138 So. 0.0004.16 2 -0.x = fH;CCOMB 2003 = 0.0095057 ka, = 10 lopka ·4·3 10 = 5.012X105 4H3(CO2H) 2002 CH3CCOLH) (600), + H+ 0.009505 m the ty ty CWe can neglect the contribution of third dissociation towards H+ concentration.